# AnyCap Project: A Unified Framework, Dataset, and Benchmark for Controllable Omni-modal Captioning

### v1

@@ Abstract
Controllable captioning is essential for precise multimodal alignment and instruction following, yet existing models often lack fine-grained control and reliable evaluation protocols. To comprehensively address this gap, we introduce the AnyCap Project, developing an integrated solution spanning model, dataset, and evaluation. Specifically, we propose AnyCapModel (ACM), a lightweight plug-and-play framework that enhances the controllability of existing foundation models for omni-modal captioning, avoiding the training of base model. ACM leverages the original caption from base models, while incorporating user instructions and modality features to produce improved captions. In addition, to make up for the data deficiencies of controllable multimodal captioning, we construct AnyCapDataset (ACD), which covers $3$ modalities, $28$ types of user instructions, and $300$k high-quality data entries. We further propose AnyCapEval, a novel benchmark that offers more reliable evaluation metrics for controllable captioning by decoupling content accuracy and stylistic fidelity. ACM significantly improves caption quality across a diverse set of base models on AnyCapEval. Notably, ACM-8B boosts GPT-4o’s content scores by $45\%$ and style scores by $12\%$. It also shows substantial gains on widely-used benchmarks such as MIA-Bench and VidCapBench.
@@ 1 Introduction
Figure 1: The overall framework of ACM. Given image, video, and audio inputs, base captioning models first generate initial outputs. ACM refines these outputs by aligning modality features with specific user instructions, enabling controllable captioning without retraining the base models.
Multimodal large language models (MLLMs) are rapidly evolving towards omni-modal intelligence, processing diverse modalities including images, videos, and audio. Textual captions serve as a universal interface linking these modalities, benefiting numerous tasks such as retrieval, question answering, and content generation [1, 2, 3, 4, 5, 6]. Recently, controllability in captioning has garnered increasing attention. It refers to the ability to generate captions that precisely follow user instructions, such as emphasizing specific aspects or adopting designated stylistic expressions. Such control is crucial for tailoring captions to user needs and enhancing downstream tasks [7, 8, 9]. Yet, achieving precise and reliable omni-modal controllable captioning remains a significant challenge, largely constrained by limitations in current framework, data, and evaluation approaches [10, 11, 12, 13, 14].
Controllable captioning across diverse modalities faces three main challenges: (1) Limited controllability capabilities of open-source models. Existing caption models typically rely on rigid control signals, such as soft prompts [13] or bounding boxes [15, 14], to adjust the fine-granularity of the caption. While this allows certain control, it limits flexibility and reduces the variety of generated results. Natural language prompts provide a more flexible way to guide captions, yet current methods often perform well only on specific aspects (e.g., detailed descriptions) due to limited and narrow training data. Retraining these models to support broader control is expensive and may weaken their overall language ability. (2) Lack of controllable omni-modal caption datasets. Manual annotation for controllable captioning is costly and labor-intensive. Meanwhile, due to the lack of low-cost and quality-assured data generation pipelines, generating large-scale data using models remains challenging. (3) Lack of appropriate evaluation benchmarks and metrics. Classic metrics (e.g., BLEU [16], CIDEr [17]) ignore content accuracy and stylistic adherence, whereas LLM-based scorers suffer from high variance, stylistic bias, and poor diagnostic power [18, 19, 20, 21, 22, 23, 24, 25].
To address these challenges, we propose AnyCapModel (ACM), a plug-and-play framework that enhances fine-grained controllability in omni-modal caption generation with frozen base models. By aligning user instructions, modality features, and initial captions, ACM produces improved captions to meet diverse control demands (Fig. 1). Without expensive training of base models, it offers a general and low-cost path toward controlled multimodal captioning.
For the lack of training data for controllable captioning, we introduce AnyCapDataset (ACD), a large-scale omni-modal dataset spanning images, videos, and audio across 28 distinct control dimensions. Each sample includes customized control signals and a pair of captions, one preferred and one rejected, clearly reflecting differences in control quality. Crucially, this preference-based design only requires that chosen caption outperform their rejected one, largely lowering annotation difficulty. This design supports learning precise alignment with diverse user preferences, which is directly compatible with our training paradigm and applicable to other preference-based methods [26, 27, 28, 29, 30, 31].
To enable a more reliable assessment for controllable captioning, we design AnyCapEval, a benchmark that assesses content and style dimensions separately. We propose Keypoint Density (KPD) and detailed scoring criteria, providing more objective and precise evaluation metrics compared to existing methods. Specifically, the benchmark evaluates both Content that focuses on adherence to control instructions, and Style that captures expressive fidelity and hallucination reduction. This two-dimensional evaluation gives rise to improved accuracy and reduced variance in assessment.
Extensive experiments demonstrate that ACM significantly improves controllability across image, video, and audio modalities, benefiting both proprietary models and various open-source alternatives. Specifically, integrating ACM-2B with GPT-4o [1] improves content controllability by 37% and style controllability by 8%. Moreover, the larger ACM-8B empowers certain open-source models, such as InternVL2.5-8B [32], to match or even surpass proprietary GPT-4o across multiple control dimensions. It consistently shows improvements on widely-used benchmarks including MIA-Bench [22] and VidCapBench [23]. In summary, our main contributions are:
(1) We introduce ACM, a versatile plug-and-play framework that boosts caption controllability for frozen base models across image, video, and audio modalities. Given the scarcity of controllable caption data, we construct ACD, a large-scale omni-modal dataset with preference pairs tailored for controllable captioning, to the best of our knowledge.
(2) We propose AnyCapEval, a novel evaluation benchmark designed for controllable captioning. It enhances assessment reliability by providing both content and style related metrics.
(3) We provide extensive experiments to show that ACM greatly improves controllability and reduces hallucinations across diverse modalities and models, achieving significant gains on both AnyCapEval and the widely-used benchmarks.
@@ 2 Related Work
Controllable captioning. Controllable captioning conditions text generation on user-instructions to dictate content and style. Early efforts in this field primarily focused on incorporating predefined and often low-level control signals. For example, researchers explored controlling caption generation based on explicit properties of the output text, such as desired caption length [11], or by conditioning on specific visual information from the image [33, 34]. This included leveraging particular visual regions or objects, sometimes specified via bounding boxes [7], or incorporating specific aspects for video captioning [35, 36]. These methods, while pioneering, lacked the expressiveness and intuitiveness of natural-language instructions that align closely with human preferences. More recent efforts leveraging MLLMs have significantly advanced image captioning, yet still often produce global and modality-specific captions, limiting their applicability to omni-modal settings [8, 9, 12, 37, 38, 39, 40, 13, 41, 42, 43, 44]. In contrast, our ACM supports flexible, language-based controls across images, videos, and audio, which can be easily extended to additional modalities.
Captioning benchmarks. Existing captioning benchmarks are inadequate for precisely evaluating caption controllability. Metrics originally developed for machine translation, such as BLEU [16] and CIDEr [17], poorly reflect semantic accuracy and controllability. Critically, their reliance on n-gram overlap fails to verify adherence to specific control signals, focusing instead on general similarity. While evaluations leveraging large language models (LLMs) can offer some insight into semantic coherence [18, 22, 45, 46, 47, 48, 49, 50, 51], they often tend to be unreliable due to significant scoring variability and potential biases [52, 53]. Similarly, evaluating adherence to fine-grained control using LLMs is often subjective and inconsistent. To address these limitations, we introduce a more fine-grained benchmark that employs keypoint density (KPD) and detailed scoring criteria, rigorously assessing controllability from both content and style dimensions.
3 AnyCap Project
In Sec. 3.1, we first propose AnyCapModel (ACM), a plug-and-play approach to help existing foundation models with caption controllability without fine-tuning. In Sec. 3.2, we present AnyCapDataset (ACD), a comprehensive dataset featuring a rich diversity of controllable captions. Finally, in Sec. 3.3, we detail the AnyCapEval benchmark for caption evaluation of control accuracy and hallucination.
3.1 AnyCapModel
Figure 2: The workflow of AnyCapModel. (a) Integrating ACM into frozen base models enhances caption controllability across multiple modalities. (b) ACM aligns modality features, initial captions, and user instructions to produce refined, instruction-compliant captions.
Workflow. As illustrated in Fig. 2(a), given a modality input $m$ from possible modalities such as image ($m^{\text{img}}$), video ($m^{\text{vid}}$), or audio ($m^{\text{aud}}$), alongside a user-specified instruction $q$, an initial caption $y_{0}$ is first generated by a frozen base model $\mathcal{M}_{g}$. Subsequently, our alignment framework ACM ($\mathcal{M}_{a}$) refines this caption:
@@ 3.1 AnyCapModel
As shown in Fig. 2(b), modality-specific encoders (e.g., InternViT [32] for $m^{\text{img}}$ and $m^{\text{vid}}$, and EAT [54] for $m^{\text{aud}}$) extract feature embeddings from respective modality inputs. These modality embeddings are projected into a shared semantic space via modality-specific linear transformations (MLPs). Concurrently, $q$ and $y_{0}$ are tokenized and embedded into textual embeddings. Finally, modality and textual embeddings are concatenated and input into an autoregressive language model to produce a refined, instruction-compliant caption $y_{c}$.
Training strategy. We adopt the residual correction training strategy [55]. This approach focuses on refining existing captions rather than generating perfectly accurate ones from scratch. To effectively train the model, ACM learns how to correct uncontrolled or hallucinated outputs by contrasting them with controlled, accurate examples. During training, we include approximately $40\%$ of data samples where initial captions already match desired controlled captions or are factually correct. This inclusion explicitly guides the model in recognizing instances requiring minimal or no correction, thereby enhancing its overall alignment capability.
Training setting. We employ the AdamW optimizer for $3$ epochs with a learning rate of $1\times 10^{-6}$, a cosine learning rate schedule ($0.03$ warmup ratio), weight decay of $0.01$, and a global batch size of $256$, using bfloat16 mixed-precision. During ACM training, the external modality backbones are frozen, while the ACM’s internal components including its language model are updated.
3.2 AnyCapDataset
Definition. This dataset is constructed to train omni-modal systems that generate descriptions under explicit user instruction. ACD adopts a carefully designed triplet structure:
$D_{\text{ACD{}}}:=\left\{(q_{i},c_{i},a_{i})\right\}_{i=1}^{N},$ (2)
where $q_{i}$ denotes an instruction that specifies the required properties of a caption. For each instance, $c_{i}$ is a high-quality caption adhering to the instruction, while $a_{i}$ is a suboptimal caption that may exhibit minor deficiencies in factual accuracy, level of detail, or compliance with the instruction.
Table 1: Supported instruction types across image, video, and audio modalities in AnyCapDataset. Checkmarks (✓) indicate where each control is applicable. Content controls regulate what is said, whereas Style controls regulate how it is articulated.
@@ 3.2 AnyCapDataset
Instance Appearance (IApp) Characterise the visual attributes of a designated instance
@@
Instruction types. The instruction types (shown in Table 1) are systematically identified by surveying caption-related literature to find common control categories and analyzing typical requirements of downstream applications. This systematic process allows us to curate a set of instruction types aiming for balanced diversity and practicality, facilitating precise control over caption generation.
Construction pipeline. For each multimodal sample, a user instruction $q$ and a related compliant caption $c$ are generated in a single step. For each instruction type, we design a detailed prompt including: (1) task description, (2) multimodal input, (3) original reference title if available, and (4) several high-quality examples that illustrate the desired content or style constraints. These prompts guide the model to produce well-aligned $(q,c)$ pairs. Further details can be found in the Appendix A.
We primarily use open-source models like InternVL2.5 [32], and apply proprietary models only for more complex cases such as perspective, balancing quality and cost. Before large-scale generation, each prompt goes through a strict validation phase: we sample about $20$ instances per modality and instruction type, and manually verify that the captions strictly follow the user instruction, contain no hallucinations, meet formatting requirements, and preserve the input’s unique features. Only prompts with a $100\%$ pass rate are used for batch generation. To generate suboptimal captions $a$, we either remove guidance or add a controlled degradation instruction (e.g., missing details, factual errors, or style violations) based on the given $(q,c)$ pair. Additionally, after data generation, a random $5\%$ sample underwent manual review, finding over $95\%$ compliance with human preferences.
Significant human and temporal resources are dedicated to establishing a scalable and cost-effective pipeline. This investment has yielded a robust dataset foundation, substantially reducing costs for subsequent research and enabling the low-cost, large-scale generation of high-quality caption data.
Statistics. Following the construction pipeline described above, ACD ultimately comprises approximately $300$k $(q,c,a)$ triplets, derived and augmented from around $75$k original samples. These original samples are collected from multiple public datasets [56, 57, 58, 59, 60, 61]. ACD spans three modalities: image ($125$k samples), video ($100$k samples), and audio ($75$k samples). The average length of optimal captions $c$ is $62$ words, while suboptimal captions $a$ average $57$ words. Further details can be found in the Appendix A.

### v2

# AnyCap: Omni-Modal Captioning with Instruction Alignment

@@ Abstract
We present AnyCap, a plug-and-play framework that brings instruction alignment to omni-modal captioning. Captions offer a unified language interface for multimodal learning, and users increasingly expect instruction-driven control over their content and style. Current caption models lack explicit instruction supervision and are weak at instruction following, while directly tuning them can degrade general language ability. Achieving instruction alignment in an omni-modal setting is harder still, as each modality calls for separate models and custom designs. To address these challenges, AnyCap leverages a residual-correction paradigm that refines uncontrolled captions from existing models to instruction-aligned ones, without re-training base models. By processing multi-modality features in a unified framework, it enables one model to serve images, videos, and audio. To address the lack of instruction-based data, we construct AnyCapData, a large-scale, high-quality corpus spanning three modalities with $28$ well-designed instruction types. For evaluation, we address the limitations of current metrics for instruction-oriented captioning by designing AnyCapEval. Its key insight is to decouple evaluation into content and style for fine-grained assessment. Extensive experiments show that on AnyCapEval and diverse public benchmarks, AnyCap consistently improves both caption quality and instruction adherence for both open-source and API-based models. Notably, AnyCap-8B boosts GPT-4o’s content scores by $46\%$ and style scores by $12\%$. Our code and models will be made publicly available.
Figure 1: We design diverse natural language instructions for controllable captioning. Given image, video, and audio inputs, base models first generate initial captions. AnyCap effectively aligns them into high-quality, instruction-aligned outputs. A single AnyCap model serves all three modalities.
@@ 1 Introduction
Multimodal large language models (MLLMs) are rapidly advancing towards omni-modal intelligence (Hurst et al., 2024; Anil et al., 2023; Xu et al., 2025; Zhao et al., 2025). Textual captions provide a natural way to connect different modalities by conveying their information in a unified language form. Beyond producing generic descriptions, users increasingly expect captions that are instruction-aligned, directly adapting to desired content and style. This can more flexibly support multimodal tasks such as retrieval (Radford et al., 2021), question answering (Li et al., 2022), and content generation (Ramesh et al., 2022).
Yet, instruction-aligned captioning faces three significant challenges. (i) Existing models have limited instruction following ability. Many are trained as general-purpose understanding models. They can lean toward the requested focus, but mainly produce broad content or unrelated details (Wang et al., 2025). For finer control over captions, current approaches usually rely on special tokens (Dwibedi et al., 2024) or bounding boxes (Zhao et al., 2024; Huang et al., 2024), to guide local descriptions. In practice, however, users could prefer flexible natural language instructions (e.g., in Table 1) to specify caption types, focuses, and styles. This represents a more fine-grained and natural form of control, yet current captioning models fall short in this ability. Bridging the gap is non-trivial, as directly fine-tuning these models to emphasize such instructions can weaken their overall language capability (Luo et al., 2023). (ii) Adapting instruction alignment to different modalities further adds complexity. Each modality imposes distinct demands. For example, visual captions often emphasize object appearance, whereas audio captions may reflect prosodic cues (Sharma et al., 2023; Mei et al., 2022). Meeting heterogeneous requirements usually forces the use of separate models. (iii) High-quality training data and benchmarks for omni-modal captioning are also lacking. Most caption datasets cover a single modality (Wang et al., 2024c; Qian et al., 2025). They also provide only coarse-grained captions, which cannot support the fine-grained instruction following required above. In particular, audio data is especially scarce, with existing corpora offering only short COCO-style captions and few long-form descriptions (Kim et al., 2019).
To tackle these challenges, we propose AnyCap, a new captioning framework characterized by two key features. It (i) generates captions that are directly controlled by user language instructions, and (ii) a single model can handle images, videos, and audio data. Our core idea, inspired by residual learning mechanism in Aligner (Ji et al., 2024), is to refine coarse captions into user-desired ones using a small model. Specifically, we leverage open-source models and available APIs to provide basic captions, and then learn how to align them with user instructions using both the modality observations and instruction texts (Fig. 1). This design simplifies the task, where the base caption supplies general content, while the learning process can focus on capturing instruction intent and refining correctness. Once trained, AnyCap can be used as a plug-and-play module to stack upon diverse captioning models, enhancing captioning without modifying their parameters, or as a flexible tool for re-annotating datasets. It is worth noting that, unlike Aligner that only corrects text modality, AnyCap significantly differs by incorporating multimodal information in a unified framework. Moreover, our aim is to achieve instruction-aligned captioning in an omni-modal setting, which is the first of this kind to our knowledge.
For the lack of instruction-based caption data, we construct a large-scale dataset termed AnyCapData. Unlike prior datasets where instruction types are fixed (e.g., “describe … in short/details” (Onoe et al., 2024)), we carefully design 28 diverse instruction dimensions spanning images, videos, and audio to better reflect real-world needs (Table 1). To support our alignment paradigm, each sample is annotated as a triplet: an instruction, a preferred caption, and a less-preferred caption, clearly capturing differences in instruction alignment quality. These triplets are generated in a data-aware manner by prompting large MLLMs with the original modality inputs, ensuring that user instructions are grounded in the source content. To further secure quality, we perform small-scale manual screening for each instruction pattern and modality before large-scale generation. With this limited human effort, we improve the quality of automated data generation.
For reliable evaluation, existing methods also remain inadequate. They typically adopt traditional machine translation scores (e.g., BLEU (Papineni et al., 2002), CIDEr (Vedantam et al., 2015)) that lack semantic awareness, or rely on direct LLM-based scoring, which suffers from high randomness across instruction types (Qian et al., 2025; Chen et al., 2025b). We thus design AnyCapEval, with the key insight to categorize diverse instruction types into two orthogonal dimensions, i.e., content and style, and evaluate them separately for focused and fine-grained evaluation. For content evaluation, we introduce the Keypoint Density (KPD) metric. It employs an automatic matcher to measure the recall of content keypoints and penalizes redundancy in excessively long captions. This encourages captions that are both precise and of appropriate length. For style evaluation, we rigorously design detailed scoring rubrics for each instruction type, combined with references to guide evaluation and reduce scoring variance. This structured design enables consistent distinction between compliant and non-compliant stylistic outputs.
Extensive experiments show that AnyCap significantly improves caption quality and instruction-following ability across image, video, and audio. It benefits both open-source and proprietary API-based models. Specifically, stacking a 2B AnyCap on GPT-4o improves content following by 35% and style controllability by 9%. The improvements are also consistently shown on widely-used benchmarks including MIA-Bench and VidCapBench, demonstrating its generalizability. We further validate AnyCap through detailed ablation studies and user studies, observing that, compared to other models, AnyCap tends to convey the instruction-required information more accurately while removing redundant content (Table 11), thereby better matching user preferences.
@@ 2 Related Work
Image, video, and audio captioning. Most early captioning models focus on a single-modality and produce only brief sentences (Vinyals et al., 2015; Venugopalan et al., 2015; Kim et al., 2019). Later research on image and video captioning moves toward longer and finer-grained descriptions (Bianco et al., 2023; Chen et al., 2024a; Ge et al., 2024; Krishna et al., 2017b). Recently, some base MLLMs have attempted to unify image and video captioning (Chen et al., 2024c; Bai et al., 2025) and even cover audio (Hurst et al., 2024; Anil et al., 2023; Xu et al., 2025). Yet, under open-ended prompts these models often drift from user-specific needs, producing irrelevant details or hallucinations that misalign with the intended caption instructions (Han et al., 2024; Huang et al., 2023). To address this, we introduce a plug-and-play model that aligns coarse captions across image, video, and audio into fine-grained, high-utility outputs.
Fine-grained captioning with instructions. A range of works use control signals as instructions to specify caption content or style. Representative approaches include employing length embedding for controlling caption length (Zeng et al., 2023), bounding boxes (Cornia et al., 2019) or masks (Wang et al., 2023) for region-level grounding, soft prompting for style or domain transfer (Zhao et al., 2024), and structured prompts to standardize instance-level descriptions (Fan et al., 2024). While effective, these control interfaces are individually designed and difficult to compose, limiting their ability to capture diverse fine-grained requirements. In contrast, we focus on flexible natural language as a unified instruction interface. Our framework aligns captions with 28 instruction types across modalities to broadly cover the spectrum of captioning needs.
Caption evaluation metrics. Traditional word-overlap metrics like BLEU (Papineni et al., 2002) and CIDEr (Vedantam et al., 2015) are widely used, but they cannot evaluate semantic consistency or instruction-following ability. This limitation arises because n-gram similarity only measures surface word overlap. Another common approach is to use LLM or MLLM as judges (Maeda et al., 2024; Qian et al., 2025; Chan et al., 2023; Sun et al., 2024; Mao et al., 2024; Dong et al., 2024; Lu et al., 2024; Liu et al., 2023; Sun et al., 2023), which offers semantic coverage but suffers from large variance and model bias (He et al., 2024; Qiu et al., 2023). Recent attempts also tailor metrics toward specificity in long captions (Chen et al., 2025a). We incorporate the strengths of these directions by introducing AnyCapEval benchmark. It explicitly separates content and style assessment with a particular focus on instruction-following evaluation. We design a new keypoint density (KPD) metric and the constructed scoring rubrics for more reliable assessment of caption quality.
3 AnyCap
3.1 Framework
Figure 2: AnyCap framework. (a) AnyCap is stacked upon various MLLMs, refining their initial captions into high-quality, instruction-aligned outputs. (b) Specifically, AnyCap takes as input the initial caption, the original modality data, and the user instruction to produce the final caption.
Given a modality input $m$ from images ($m^{\text{img}}$), videos ($m^{\text{vid}}$), or audio ($m^{\text{aud}}$), captioning across different modalities often relies on separate MLLMs, and the resulting captions often fail to follow user-specified content or style requirements. To address this, we design AnyCap as a plug-and-play approach to help existing foundation models align captions with user instructions. As illustrated in Fig. 2(a), a frozen base model first processes $m$ to produce an initial caption $y_{0}$. Then, AnyCap ($\mathcal{M}_{a}$) takes the original input $m$, the initial caption $y_{0}$, and a detailed user instruction $q$ to generate a refined caption $y_{c}$, formulated as
@@ 3.1 Framework
Concretely, as shown in Fig. 2(b), AnyCap first employs modality-specific encoders (e.g., InternViT (Chen et al., 2024c) for $m^{\text{img}}$ and $m^{\text{vid}}$, and EAT (Chen et al., 2024b) for $m^{\text{aud}}$) to extract features from each modality input. These modality features are projected into a shared semantic space via modality-specific linear transformations (MLPs). Concurrently, $q$ and $y_{0}$ are tokenized and embedded into textual embeddings. Finally, modality and textual embeddings are concatenated and fed into a large language model to produce the refined, instruction-compliant caption $y_{c}$.
The architecture of AnyCap allows us to adopt a residual-correction strategy (Ji et al., 2024). Although supervision still follows standard next-token prediction on the final caption $y_{c}$, AnyCap focuses on learning the correction pattern from the initial caption $y_{0}$ to the instruction-compliant target. This refinement is also conditioned on both the modality features and the user instruction, ensuring that the final caption remains grounded in the source input while aligning with the specified requirements. With $y_{0}$ as reference, we observe that residual learning makes the task easier and delivers two gains: (i) it outperforms directly fine-tuning base MLLMs with instructions alone, as shown in Secs. 4.3 and 9; and (ii) it enables a single model to be trained jointly across image, video, and audio data. We only train the lightweight plug-and-play AnyCap, leaving the base MLLMs frozen to preserve its original capabilities. To help the model recognize when no change is needed, about 40% of training instances already satisfy the instruction. It makes training more stable and mitigates over-editing issue, as ablated in Fig. 5.
3.2 AnyCapData
We propose AnyCapData that features three aspects: (i) it covers large-scale, high-quality caption data across images, videos, and audio; (ii) each caption is paired with user instructions to specify requirements; and (iii) each sample adopts a triplet structure to support our residual-correction training, formulated as $(q,c,a)$. Here, $q$ denotes the language instruction, $c$ is a high-quality caption adhering to the instruction, while $a$ is a suboptimal caption that may exhibit minor deficiencies in factual accuracy, level of detail, or compliance with the instruction.
Table 1: Supported instruction types across image, video, and audio modalities in AnyCapData. Checkmarks (✓) indicate where each control is applicable. Content types control what information is conveyed, while Style types shape how the message is delivered.
@@ 3.2 AnyCapData
Instance Appearance (IApp) Characterise visual attributes of a designated instance
@@
Instruction types. To ensure both diversity and practicality, we identify instruction types by systematically surveying caption-related literature to surface common control categories and by analyzing typical downstream requirements. The final selections are further conditioned on the distinctive characteristics of each modality, as listed in Table 1. More details are illustrated in Sec. A.2.
Construction pipeline. With the designed instruction types, we first use MLLMs to jointly generate a specific user instruction $q$ and its compliant caption $c$ for each image, video, or audio sample. This avoids cases where a separately predicted instruction might request information the caption cannot provide. The prompts for data construction are tailored per modality and instruction type, including clear task descriptions, explicit length and style constraints, common pitfalls, few-shot exemplars, and the original media as input. For audio, we include the dataset’s reference caption (when available) in the prompt to reduce hallucination. Importantly, $q$ is instantiated for each example (e.g., “describe the watermelon’s appearance in this image”, rather than a generic “describe details”), enabling fine-grained and targeted control. We primarily adopt strong open-source VLMs (e.g., InternVL2.5-78B, Qwen2.5-VL-72B) for visual modality. GPT-4o are selected for harder controls (e.g., perspective, poetry) and for audio modality, balancing coverage and cost. For contrastive supervision, given $(q,c)$ we derive a suboptimal caption $a$ by removing guidance to make it uncontrolled or injecting degradations to increase hallucination. The latter includes omitting required keypoints, introducing mild factual errors, or violating style/length constraints.
Before large-scale automatic construction, each prompt template is manually validated on about 20 instances per modality–type. The validation requires full instruction compliance, no hallucination, correct formatting, and preservation of modality-specific details. After large-scale generation, a check of 5% of the data shows over 95% agreement with human preferences. We also apply type-specific length gate and automatic sanity check to filter degenerate or repetitive outputs. The entire pipeline requires only limited human intervention to enable low-cost scaling.
Statistics. The resulting AnyCapData contains about 300300k $(q,c,a)$ triplets, spanning images (125k), videos (100k), and audio (75k), constructed from and augmented over roughly 75k samples, which are derived and augmented from 7575k original samples. These original samples are collected from multiple public datasets (Wang et al., 2024b; Onoe et al., 2024; Urbanek et al., 2024; Cui et al., 2024; Fan et al., 2024; Kim et al., 2019) with the ensured quality. Data sources and splits, detailed construction procedures, and statistic details are shown in Appendix A.